OpenStudy (anonymous):

Solve the equation for x: lnx^1/2-2ln3=0

OpenStudy (anonymous):

So far I got ln (x^1/2 / 2.3)=0 I JUST DON'T KNOW HOW TO SOLVE FOR X FROM HERE

OpenStudy (mertsj):

\[\ln \frac{x ^\frac{1}{2}}{9}=0\]

OpenStudy (anonymous):

\[\ln (x)^{1/2}-2 \ln (3) = 0 \rightarrow \frac{ 1 }{ 2 } \ln (x) - 2 \ln(3) = 0 \rightarrow \ln (x) = 4 \ln (3) \rightarrow \ln(x) = \ln(3)^4\]\[\rightarrow \ln(x) = \ln(81) \implies x = 81\] @aebiri2

OpenStudy (mertsj):

\[e^0=\frac{\sqrt{x}}{9}=1\]

OpenStudy (mertsj):

Square both sides: \[\frac{x}{81}=1\]

OpenStudy (mertsj):

\[x=81\]

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

I understood both methods